What is the distance between the foci of an ellipse when the lengths of the major and minor axes are 52 and 20?
Answer by lwsshak3(11628)

What is the distance between the foci of an ellipse when the lengths of the major and minor axes are 52 and 20?
***
given length of major axis=52=2a
a=26
a^2=676
given length of minor axis=20=2b
b=10
b^2=100
c^2=a^2-b^2=676-100=576
c=√576=24(distance from center to foci)
2c=48=distance between the foci